OpenStudy (anonymous):

1. Suppose that y varies directly with x and inversely with z, and y = 18 when x = 15 and z = 5. Write the equation that models the relationship. Then find y when x = 21 and z = 7. (1 point) y = 5z over x; five-thirds t = 5x over z; 15 t = 6x over z; 18 t = 6z over x; 2

OpenStudy (anonymous):

I would be so helpful if I was to know how to plug it in

OpenStudy (anonymous):

y=k x/z y = 18 when x = 15 and z = 5 pluuging these values we have 18=k(15)/5 or k=6 hence y=6x/z now y =6(21)/7 =6*3=18 hence y= 6x over z; 18

OpenStudy (anonymous):

So it doesn't matter if it is a t= or a y= ?

OpenStudy (anonymous):

it should be y there might be a misprint or typo error...

OpenStudy (anonymous):

probably so...but thank you so much!!!! I really appreciate it( :